OpenStudy (anonymous):

What are the intercepts for the line 3x + 4y = 7?

OpenStudy (pradipgr817):

make it in the form y=mx+c c is the intercept

hartnn (hartnn):

to get x-intercept, put y=0 to get y-intercept, put x=0

OpenStudy (pradipgr817):

if you want intercept on both axis, put x=0, y intercept put y=0, x intercept

OpenStudy (anonymous):

3x + 4y = 7 3x + 4y -7=0 y intercept = -c/b

OpenStudy (anonymous):

divide the whole equation with rhs on both sides... then 3x/7 + 4y/7 =1 then bring the numerator down... x/(7/3) + y/(7/4) =1 now we have our intercepts, a= 7/3 , b= 7/4